# Identifying Best Candidates for Busbar Splitting

### v1

@@ 2 Preliminaries: Topology Optimization with Busbar Splitting
(M1.2) and (M1.3) are related to the voltage angles and magnitudes of the two buses at the extremes of each ($sw^{ac}_{\upsilon mi}$) switching element, with $M_{\theta},M_{m}$ being big-enough M values ($2\pi,1$). (M1.4) and (M1.5) limit the active and reactive powers of all switching elements to its maximum $\overline{P}^{sw,ac}_{\upsilon mi}$, $\overline{Q}^{sw,ac}_{\upsilon mi}$ and minimum values $\underline{P}^{sw,ac}_{\upsilon mi}$, $\underline{Q}^{sw,ac}_{\upsilon mi}$. (M1.6) and (M1.7) are “exclusivity” constraints. Line switching is allowed to be performed by modeling (M1.6) as an inequality ($\leq$ 1) constraint. Therefore, switches $z^{sw,ac}_{\upsilon mi}$ and $z^{sw,ac}_{\kappa mi^{\prime}}$, connecting each network element to the parts of the split busbar, are allowed to be open simultaneously. Note that (M1.7) imposes that if the busbar coupler is closed, the switch $sw^{ac}_{\kappa mi^{\prime},t}$ connecting the element to the original busbar is always closed.
Besides the mixed-integer constraints describing the BuS logic, the LPAC-BuS model includes more standard OPF constraints. These are all linear except for the quadratic constraint describing the LPAC approximation term $cs_{i,j}$ of the cosine of of the angle difference $cos(\theta_{i}-\theta_{j})$ between the end buses $i$ and $j$ of each branch $lij\in\mathcal{T}^{ac}\cup\mathcal{T}^{ac,rev}$ [38].
3 Proposed Metrics to Select Candidate Busbars for Busbar Splitting
@@ 3.1 Metric design
When designing these metrics, we only use information that can be obtained with comparatively low computational effort, such as from direct analysis of the system topology,e.g. identifying the number of elements connected to a bus, or from solving an AC-OPF. We further take inspiration from metrics previously proposed to identify candidate busbars for splitting in BuS. Specifically, we consider the general idea of identifying buses that are associated with congestion, have large differences in LMP to neighboring buses, and where a significant number of elements are connected to the busbar. Based on this, we propose the following three metrics:
3.1.1 Metric $\phi$ – Sum of absolute LMP differences
@@
This metric is defined as the sum of the absolute differences in LMPs $|\Delta_{LMP}|$ across all branches connected to the selected busbar. This metric rewards buses that have multiple, high $\Delta_{LMP}$ values and is inspired by congestion zone identification, $\Delta_{LMP}$, and # of elements connected to the selected busbar metrics, If a busbar lies within a congestion zone, even if one of its branches exhibits a large $\Delta_{LMP}$, the remaining connections are likely to have limited $\Delta_{LMP}$ values, thus limiting the value of BuS to reduce generation costs [39]. Moreover, the potential to find topological actions reducing the total costs increases with the number of elements connected to a busbar [22, 23]. Thus, busbars exhibiting large $\Delta_{LMP}$ values over multiple incident branches are promising candidates.
3.1.2 Metric $\xi$ – Number of congested branches
@@
This metric evaluates the loading of the branches incident to a busbar and identifies the number of congested branches, inspired by a combination of the congestion zones identification and # of elements connected to the busbar metrics. A significant disparity in utilization of branches connected to the same busbar indicates the presence of local bottlenecks, where one or more branches operate closer to their thermal limits while others remain underutilized. Such imbalances suggest opportunities for remedial actions to redistribute flows, alleviate congestion, and therefore reduce generation costs.
3.1.3 Metric $\zeta$ – Binding voltage angle and magnitude limits
@@
While a majority of existing work on OTS and BuS metrics has only considered DC-OPF representations, we base our metrics and optimization on AC-OPF (and approximations thereof) that also includes consideration of reactive power and voltage magnitudes. Specifically, the OPF includes voltage magnitude, voltage difference and angle difference constraints for each branch. Branches with binding voltage magnitude or angle difference limits imply congestion and restricted power transfer capability. Identifying busbars connected by branches close to these operational limits highlights locations where topological actions can increase the grid transmission capacity.
3.2 Evaluation, testing and validation
@@
To evaluate the metrics, we first solve an AC-OPF with the original topology to determine the generation costs. Then, we solve the LPAC-BuS model with one busbar selected for topology optimization, and finally run an AC-OPF with the optimized topology to evaluate the cost reduction.
@@
4 Results and discussion
4.1 Hardware setup
Table 2: Test cases from pglib [40] used in this work to test and validate the proposed metrics.
@@ 4.1 Hardware setup
The results of the OPF and LPAC-BS models are computed using a MacBook Pro with chip M1 Max and 32 GB of memory, using Gurobi v12.0, MIP gap 0.01% for LPAC-BS, and Ipopt with linear solver ma97 for the OPF model.
4.2 Testing Results
We perform the testing phase using the 39-bus and 188-bus test cases introduced in Table 2.
4.2.1 Accuracy and Computational Efficiency
Using the 39-bus test case as an example, we first make some basic observations about the accuracy and computational effort associated with our approach. Fig. 2 plots the $\phi$-metric of each busbar(representing the sum of absolute LMP differences) against the reduction in generation cost achieved by splitting that bus. Each busbaris represented by two dots, where the blue dot represent the results achieved with an LPAC-OPF feasibility check and the green dot represents the results achieved with the AC-OPF feasibility check. We observe that there are some differences between the two models, but that the trends (e.g. busbars with highest $\phi$ values and highest generation cost reductions) are similar. This shows that the LPAC-BuS has sufficient accuracy to be used for topology optimization.
Regarding computational efficiency, the whole optimization process (splitting one busbar per time with LPAC-BuS, plus AC- and LPAC-FC) took 11.4 s, confirming the efficiency of the proposed LPAC-BuS model.
4.2.2 Testing of Metrics on 39-bus test case
@@
Further analysis of the results yields some interesting insights. First, the branch connecting busbars 2 and 30 has the highest LMP difference $\Delta_{LMP}$. However, when either of these buses are allowed to be split, the optimal topology is the same as the original one. Thus, there is no change in generation cost. This finding confirms that the widely used metric $\Delta_{LMP}$ is insufficient, and that we need more (or combined) metrics to select relevant busbars for topology optimization. Further, busbar 30 has only 2 elements connected to it and its voltage magnitude $U^{m}_{30}$ is at the upper limit 1.1. This suggests that it is important to also consider the # of elements connected to the selected busbar and the $\zeta$-metric representing the presence of binding voltage limits. Busbars 2 has 2 congested lines (used almost at full capacity) connected to it, confirming that the $\xi$-metric representing the number of congested lines might be important.
Figure 2: For each bus in 39-bus test case, the optimal topology from the LPAC-BuS model is tested for an (“AC”- and “LPAC”-) OPF feasibility check (FC). The differences in total costs of the FC compared to the original OPF are plotted on the x-axis, while the y-axis orders the busbar according to each bus’ sum of the absolute differences in LMPs over the branches connected to it ($\phi$). The LPAC-FC and AC-FC see reductions in costs for the same buses.
4.2.3 118-bus test case
We apply the previous findings to select busbars for splitting in the 118-bus case. Table 3 shows the value of the different metrics as well as the cost reduction achieved with BuS for the buses with the highest $\phi$-metric values (top) and busbars that lead to a cost decrease of >>0.05%, but are not in the top-10 $\phi$-ranked busbars (bottom).
We observe that the busbars with high $\phi$, many connected several branches and without binding voltage limits ($\zeta$) lead to a cost decrease compared to the AC-OPF. Importantly, 4 out of 5 busbars with a cost reduction larger than 0.1% are included among the busbars with the highest $\phi$ values. The highest cost decrease is for busbar 69, whose optimized topology is shown in Fig. 3.
@@
Among the buses in the bottom part of Table 3, we observe that busbars 80 and 42 are connected to several branches. The other busbars decrease cost by not reconnecting branches to the split busbar, essentially performing transmission line switching rather than BuS. They are therefore more complex to identify with our metrics. Note that all the busbars with a cost decrease of >>0.05% are within the first half of the 118 $\phi$-ranked busbars, and two of them have $\phi$ above its average value of 219.13 [\$/MWh]. Note that for this test case, the entire optimization process took 214.7 s. Based on these results, we consider busbars with several connected branches/elements and a $\phi$ above average in the validation step.
Table 3: Metrics of the 10 busbars in the 118-bus test case with the highest sum of the absolute differences in LMPs over their branches ($\phi$), and of the busbars with the highest cost decrease wrt AC-OPF, not in the top 10 $\phi$-ranked busbars. Busbars at voltage limit ($\zeta$) do not have a cost decrease, while there are limited congested (>> 85% utilization) branches ($\xi$).
@@
4.3 Validation Results
In this step, we apply the metrics discussed above to identify busbars for splitting in the 793-bus and 3374-bus cases listed in Table 2.
4.3.1 793-bus test case
We first select the 15 busbars with the highest $\phi$-values, and remove the busbars with fewer than 4 elements connected to them. As a result, the 8 busbars on the top part of Table 4 are selected. Second, we select the 7 busbars with the highest # of elements and branches connected to them, and have a $\phi$ higher than the mean $\phi$ value. These busbars are displayed on the bottom part of Table 4. We ensure that the selected busbars are connected to at most one congested line ($\xi$-metric $\leq 1$) and have no binding voltage angle or magnitudes constraints ($\zeta$-metric =0=0).
Table 4: Identification of 15 candidates for busbar splitting in 793-bus test case with the proposed metrics, and related cost decrease. 12 optimized topologies out of 15 lead to a cost decrease in the AC-OPF.

### v2

@@ II-B Grid topology optimization model
Moreover, the reasoning about BuS introduced in the previous Section is formulated through constraints (M1.12) to (M1.18) and applied to a set of switches $\mathcal{T}^{\text{sw,ac}}$ and busbar couplers $\mathcal{T}^{\text{ZIL,ac}}$, with each switching element represented by a binary variable. (M1.13) and (M1.14) are related to the voltage angles and magnitudes of the two buses at the extremes of each ($sw^{ac}_{\upsilon mi}$) switching element, with $M_{\theta},M_{m}$ being big-enough M values ($2\pi,1$). (M1.15) and (M1.16) limit the active and reactive powers of all switching elements to its maximum $\overline{P}^{sw,ac}_{\upsilon mi}$, $\overline{Q}^{sw,ac}_{\upsilon mi}$ and minimum values $\underline{P}^{sw,ac}_{\upsilon mi}$, $\underline{Q}^{sw,ac}_{\upsilon mi}$. (M1.7) and (M1.8) are “exclusivity” constraints. Line switching is allowed to be performed by modeling (M1.7) as an inequality ($\leq$ 1) constraint. As a result, switches $z^{sw,ac}_{\upsilon mi}$ and $z^{sw,ac}_{\kappa mi^{\prime}}$, connecting each network element to the parts of the split busbar, are allowed to be open simultaneously, therefore removing the network element from the split busbar. Note that (M1.7) imposes that if the busbar coupler is closed, the switch $sw^{ac}_{\kappa mi^{\prime},t}$ connecting the element to the original busbar is always closed.
Model 1: Grid topology optimization model, LPAC formulation
Minimize:
$\sum_{k\in G}c_{2k}\cdot{S^{g}_{k}}^{2}+c_{1k}\cdot S^{g}_{k}+c_{0k}+\sum_{ZILii^{\prime}\in\mathcal{T}^{\text{ZIL,ac}}}c_{ZIL}\cdot z^{sw,ac}_{ZILii^{\prime}}$ (M1.1)
Subject to:
AC bus:
$\theta_{r}=0$ (M1.2)
$(\underline{U}^{m}_{i}-1)\leq\phi_{i}\leq(\overline{U}^{m}_{i}-1)\quad\forall i\in\mathcal{I},$ (M1.3)
$\underline{\theta}_{i}\leq\theta_{i}\leq\overline{\theta}_{i}\quad\forall i\in\mathcal{I},$ (M1.4)
$\sum_{\begin{subarray}{c}k\in\mathcal{G}_{i}\end{subarray}}S^{g}_{k}+\sum_{\begin{subarray}{c}l\in\mathcal{M}_{i}\end{subarray}}S^{m}_{l}-Y^{s}_{i}(1+2\phi_{i})+\sum_{\begin{subarray}{c}\upsilon mi\in(\mathcal{T}^{\text{sw,ac}}_{i}\cup\mathcal{T}^{\text{ZIL,ac}}_{i})\end{subarray}}{S^{sw,ac}_{\upsilon mi}}=\sum_{\begin{subarray}{c}lij\in\mathcal{T}^{ac}\end{subarray}}S^{ac}_{lij}$ $\forall i\in\mathcal{I},$ (M1.5)
Generator
$\underline{S}^{g}_{k}\leq S^{g}_{k}\leq\overline{S}^{g}_{k},\quad\forall k\in\mathcal{G}$ (M1.6)
Branches
$P_{lij}=(g_{s,ij}+g_{ij})(1+2\cdot\phi_{i})-g_{ij}(cs_{ij,wt}+\phi_{i}+\phi_{j})-b_{ij}(\theta_{i}-\theta_{j})\quad\forall lij\in\mathcal{T}^{ac}\cup\mathcal{T}^{ac,rev}$ (M1.7)
$Q_{lij}=(b_{s,ij}+b_{ij})(1+2\cdot\phi_{i})-b_{ij}(cs_{ij,wt}+\phi_{i}+\phi_{j})-g_{ij}(\theta_{i}-\theta_{j})\quad\forall lij\in\mathcal{T}^{ac}\cup\mathcal{T}^{ac,rev}$ (M1.8)
$|P_{lij}|\leq\overline{P}_{lij}\quad\forall lij\in\mathcal{T}^{ac}\cup\mathcal{T}^{ac,rev}$ (M1.9)
$|Q_{lij}|\leq\overline{Q}_{lij}\quad\forall lij\in\mathcal{T}^{ac}\cup\mathcal{T}^{ac,rev}$ (M1.10)
$cs_{ij,wt}\leq 1-\frac{(1-cos(\overline{\Delta\theta_{ij}})}{(\overline{\Delta\theta_{ij}})^{2}}(\theta_{i}-\theta_{j})^{2}\quad\forall lij\in\mathcal{T}^{ac}\cup\mathcal{T}^{ac,rev}$ (M1.11)
Busbar Splitting
$(1-z^{sw,ac}_{\upsilon mi})\cdot M_{\theta}\leq\theta_{m}-\theta_{i}\leq(1-z^{sw,ac}_{\upsilon mi})\cdot M_{\theta},\quad\forall\upsilon mi\in\mathcal{T}^{\text{sw,ac}}\cup\mathcal{T}^{\text{ZIL,ac}}$ (M1.12)
$(1-z^{sw,ac}_{\upsilon mi})\cdot M_{m}\leq\phi_{m}-\phi_{i}\leq(1-z^{sw,ac}_{\upsilon mi})\cdot M_{m},\quad\forall\upsilon mi\in\mathcal{T}^{\text{sw,ac}}\cup\mathcal{T}^{\text{ZIL,ac}}$ (M1.13)
$z^{sw,ac}_{\upsilon mi}\cdot\underline{P}^{sw,ac}_{\upsilon mi}\leq P^{sw,ac}_{\upsilon mi}\leq z^{sw,ac}_{\upsilon mi}\cdot\overline{P}^{sw,ac}_{\upsilon mi},\quad\forall\upsilon mi\in\mathcal{T}^{\text{sw,ac}}\cup\mathcal{T}^{\text{ZIL,ac}}$ (M1.14)
$z^{sw,ac}_{\upsilon mi}\cdot\underline{Q}^{sw,ac}_{\upsilon mi}\leq Q^{sw,ac}_{\upsilon mi}\leq z^{sw,ac}_{\upsilon mi}\cdot\overline{Q}^{sw,ac}_{\upsilon mi},\quad\forall\upsilon mi\in\mathcal{T}^{\text{sw,ac}}\cup\mathcal{T}^{\text{ZIL,ac}}$ (M1.15)
$({P^{sw,ac}_{\upsilon mi}})^{2}+({Q^{sw,ac}_{\upsilon mi}})^{2}\leq z^{sw,ac}_{\upsilon mi}\cdot({\overline{S}^{sw,ac}_{\upsilon mi}})^{2},\quad\forall\upsilon mi\in\mathcal{T}^{\text{sw,ac}}\cup\mathcal{T}^{\text{ZIL,ac}},$ (M1.16)
$z^{sw,ac}_{\upsilon mi}+z^{sw,ac}_{\kappa mi^{\prime}}\leq 1,\quad\forall(\upsilon mi,\kappa mi^{\prime})\in\mathcal{T}^{\text{sw,ac}}$ (M1.17)
$z^{sw,ac}_{\kappa mi^{\prime}}\leq(1-z^{sw,ac}_{ZILii^{\prime}}),\quad\forall(\kappa mi^{\prime},ZILii^{\prime})\in\mathcal{T}^{\text{sw,ac}}\cup\mathcal{T}^{\text{ZIL,ac}}$ (M1.18)
III Proposed Metrics to Select Candidate Busbars for Busbar Splitting
@@ III-A Metric design
When designing these metrics, we only use information that can be obtained with comparatively low computational effort, such as from direct analysis of the system topology,e.g., identifying the number of network elements connected to a busbar, or from solving an AC-OPF. We further take inspiration from the metrics previously discussed in Table I to identify candidate lines to be switched in OTS problems and busbars for splitting in BuS problems. Specifically, we consider the general idea of identifying buses that are associated with congestion, have large differences in LMP to neighboring buses, and where a significant number of elements are connected to the busbar. Based on this, we propose the following three metrics:
III-A1 Metric $\phi$ – Sum of absolute LMP differences
@@
This metric is defined as the sum of the absolute differences in LMPs $|\Delta_{LMP}|$ across all branches connected to the selected busbar. This metric rewards buses that have multiple, high $\Delta_{LMP}$ values and is inspired by congestion zone identification, $\Delta_{LMP}$, and # of elements connected to the selected busbar metrics. If a busbar lies within a congestion zone, even if one of its branches exhibits a large $\Delta_{LMP}$, the remaining connections are likely to have limited $\Delta_{LMP}$ values, thus limiting the value of BuS to reduce generation costs [9]. Moreover, the potential to find topological actions reducing the total costs increases with the number of elements connected to a busbar [19, 34]. Thus, busbars exhibiting large $\Delta_{LMP}$ values over multiple incident branches are promising candidates.
III-A2 Metric $\xi$ – Number of congested branches
@@
This metric evaluates the loading of the branches incident to a busbar and identifies the number of congested branches, inspired by a combination of the congestion zones identification and # of elements connected to the busbar metrics. A significant disparity in utilization of branches connected to the same busbar indicates the presence of local bottlenecks, where one or more branches operate closer to their thermal limits while others remain underutilized. Such imbalances suggest opportunities for remedial actions to redistribute flows, alleviate congestion, and therefore reduce generation costs by trying to use the cheapest generators to their maximum capacity.
III-A3 Metric $\zeta$ – Binding voltage angle and magnitude limits
@@
While a majority of existing work on OTS and BuS metrics has only considered DC-OPF representations, we base our metrics and optimization on AC-OPF (and approximations thereof) that also includes consideration of reactive power and voltage magnitudes. Specifically, the OPF includes voltage magnitude, voltage difference, and angle difference constraints for each branch. Branches with binding voltage magnitude or angle difference limits imply congestion and restricted power transfer capability. Identifying busbars connected by branches close to these operational limits highlights locations where topological actions can increase (or not) the grid transmission capacity.
III-B Evaluation, testing and validation
@@
To evaluate the metrics, we first solve an AC-OPF with the original topology to determine the generation costs. Then, we solve the LPAC-BuS model with one busbar selected for each run of the topology optimization model, and finally run an AC-OPF with the optimized topology as a feasibility check and to evaluate the cost reduction.
@@
Table III lists the test cases from pglib [2] used in the testing and validation steps, including the number of generators, buses, branches, and loads in each test case. They range from 39 to 3374 buses.
TABLE III: Test cases from pglib [2] used in this work to test and validate the proposed metrics.
@@
III-C Simulations setup
The results of the OPF and LPAC-BS models are computed using a MacBook Pro with chip M1 Max and 32 GB of memory, using Gurobi v12.0, MIP gap 0.01%, maximum simulation time for LPAC-BS, and Ipopt with linear solver ma97 for the OPF model.
IV Results and discussion
IV-A Testing phase results
We perform the testing phase using the 39-bus and 188-bus test cases introduced in Table III.
IV-A1 Accuracy and Computational Efficiency
Using the 39-bus test case as an example, we first make some basic observations about the accuracy and computational effort associated with our approach. Fig. 2 plots the $\phi$-metric of each busbar (representing the sum of absolute LMP differences across all the branches connected to the selected busbar) against the reduction in generation cost achieved by splitting that bus. Each busbar is represented by two dots, where the blue dot represents the results achieved with an LPAC-OPF feasibility check (FC) and the green dot represents the results achieved with the AC-OPF FC. The FCs are run on the optimized topology computed through the LPAC-BuS model.
We observe that there are some differences between the two models, but that the trends (e.g. busbars with the highest $\phi$ values and the highest generation cost reductions) are similar. This shows that topology optimization through the LPAC-BuS model provides results that are sufficiently accurate to reduce the total costs of an AC-FC compared to the AC-OPF of the original topology. Therefore, the LPAC-BuS model computes topologies that are optimal and lead to a reduction in total costs compared to the original topology, even for AC-OPF formulations.
Regarding computational efficiency, the whole optimization process (splitting one busbar per time with LPAC-BuS, plus AC-FC and LPAC-FC) took 11.4 s, confirming the efficiency of the proposed LPAC-BuS model.
IV-A2 Testing of Metrics on 39-bus test case
@@
Further analysis of the results yields some interesting insights. First, the branch connecting busbars 2 and 30 has the highest LMP difference $\Delta_{LMP}$. However, when either of these buses is allowed to be split, the optimal topology is the same as the original one. Thus, there is no change in generation cost. This finding confirms that the widely used metric $\Delta_{LMP}$ is insufficient, and that we need more (or combined) metrics to select relevant busbars for topology optimization. Further, busbar 30 has only 2 elements connected to it, and its voltage magnitude $U^{m}_{30}$ is at the upper limit 1.1. This suggests that it is important to also consider the # of elements connected to the selected busbar and the $\zeta$-metric representing the presence of binding voltage angle and magnitude limits. Busbars 2 has 2 congested lines (used almost at full capacity) connected to it, confirming that the $\xi$-metric representing the number of congested lines might be important, but not when the busbar resides in a congested zone, as we commented in Section III-A.
Figure 2: For each bus in 39-bus test case, the optimal topology from the LPAC-BuS model is tested for an “AC”-OPF and “LPAC”-OPF feasibility check (FC). The differences in total costs of the FC compared to the original OPF are plotted on the x-axis, while the y-axis orders the busbar according to each bus’ sum of the absolute differences in LMPs over the branches connected to it ($\phi$). The LPAC-FC and AC-FC see reductions in costs for the same buses.
IV-A3 118-bus test case
We apply the previous findings to select busbars for splitting in the 118-bus case. Table IV shows the value of the different metrics as well as the cost reduction achieved with BuS for the buses with the highest $\phi$-metric values (top) and busbars that lead to a cost decrease of >>0.05%, but are not in the top-10 $\phi$-ranked busbars (bottom).
We observe that the busbars with high $\phi$, many connected branches/elements and without binding voltage limits ($\zeta$) lead to a cost decrease compared to the AC-OPF. Importantly, 4 out of 5 busbars with a cost reduction larger than 0.1% are included among the busbars with the highest $\phi$ values. The highest cost decrease is for busbar 69, whose optimized topology is shown in Fig. 3. In this optimized topology, power can directly flow from busbar 49 to 47, as the branches connecting them are decoupled from the original busbar 69.
@@
Among the buses in the bottom part of Table IV, we observe that busbars 80 and 42 are connected to several branches. The other busbars decrease cost by not reconnecting branches to the split busbar, essentially performing transmission line switching rather than BuS. They are therefore more complex to identify with our metrics. Finally, all the busbars with a cost decrease of >>0.05% are within the first half of the 118 $\phi$-ranked busbars, and two of them have $\phi$ above its average value of 219.13 [\$/MWh].
Note that for this test case, the entire optimization process took 214.7 s. Based on these results, we consider busbars with several connected branches/elements and a $\phi$ above average as metrics to select relevant busbars in the validation step.
TABLE IV: Metrics of the 10 busbars in the 118-bus test case with the highest sum of the absolute differences in LMPs over their branches ($\phi$), and of the busbars with the highest cost decrease wrt AC-OPF, not in the top 10 $\phi$-ranked busbars. Busbars at voltage limit ($\zeta$) do not have a cost decrease, while there are limited congested (>> 85% utilization) branches ($\xi$).
@@
IV-B Validation phase results
In this step, we apply the metrics discussed above to identify busbars for splitting in the 793-bus and 3374-bus cases listed in Table III.
IV-B1 793-bus test case
We first select the 15 busbars with the highest $\phi$-values and remove those with fewer than 4 elements connected to them. As a result, the 8 busbars on the top part of Table V are selected. Second, we select the 7 busbars with the highest # of elements and branches connected to them, and have a $\phi$ higher than the mean $\phi$ value. These busbars are displayed on the bottom part of Table V. We ensure that all the selected busbars are connected to at most one congested line ($\xi$-metric $\leq 1$) and have no binding voltage angle or magnitudes constraints ($\zeta$-metric =0=0).
TABLE V: Identification of 15 candidates for busbar splitting in the 793-bus test case with the proposed metrics, and related cost decrease. 12 optimized topologies out of 15 lead to a cost decrease in the AC-OPF.